OpenStudy (anonymous):

What is the solution to the rational equation 3x/4 + 1/4 = 5 ?

OpenStudy (mertsj):

Multiply both sides by 4 to clear the fractions.

OpenStudy (mertsj):

\[\frac{3x}{4}\times 4 +\frac{1}{4}\times 4=5\times 4\]

OpenStudy (mertsj):

\[3x+1=20\]

OpenStudy (mertsj):

Finish it.